# Improving Recursive Transformers with Mixture of LoRAs

### v1

@@ Abstract
Parameter sharing in recursive transformers reduces model size but collapses layer-wise expressivity. We propose Mixture of LoRAs (MoL), a lightweight conditional-computation mechanism that inserts Low-Rank Adaptation (LoRA) experts inside a shared feed-forward network (FFN). MoL enables token-conditional weight-space modulation of the shared FFN without untying backbone parameters, unlike prior approaches that add fixed or externally attached adapters. We pretrain a modernised recursive architecture, ModernALBERT, integrating rotary embeddings, GeGLU, FlashAttention, and a distillation-based initialisation. Across GLUE, SQuAD-v2, and BEIR, ModernALBERT (50M–120M) achieves state-of-the-art performance among compact models and surpasses larger fully parameterised baselines. We also propose an expert-merging procedure that compresses MoL into a single adapter at inference while preserving accuracy, enabling efficient deployment. Our results show that conditional weight-space modulation effectively restores the expressivity lost under aggressive parameter sharing in recursive transformers.
@@ 1 Introduction
Existing approaches attempt to mitigate the expressivity loss caused by parameter sharing in three main ways. (i) Static adapters (e.g., bottleneck adapters or LoRAs) add depth-specific parameters but do not adapt at the token level and do not modify the shared FFN itself. (ii) Mixture-of-Adapters (MoA) and LoRA-based MoE variants introduce dynamic experts after the FFN, leaving its internal transformation unchanged. (iii) Relaxed Recursive Transformers (RRT) attach static LoRA modules to a shared block, partially restoring depth-specific behaviour but lacking token-dependent routing. None of these designs introduce conditional computation inside the shared FFN during pretraining, nor do they recover layer-wise functional diversity without untying parameters across depth.
To address these limitations, we introduce ModernALBERT, a recursive transformer that restores expressivity through Mixture of LoRAs (MoL), a mechanism for conditional weight-space modulation inside the shared FFN. MoL embeds a small pool of LoRA experts directly into the FFN weights and routes tokens to a sparse subset of them, allowing each token to induce distinct low-rank updates while the backbone remains fully shared. By modulating the shared FFN itself, rather than attaching external adapters, MoL reestablishes the layer-wise functional diversity that parameter sharing typically eliminates, all while preserving strict parameter efficiency (hu2021lora).
@@
  • We introduce Mixture of LoRAs (MoL), a conditional-computation layer that applies low-rank expert updates inside a shared FFN, restoring expressivity in recursive transformers.
@@ 3.1 Parameter-Sharing through Recursions
$|\Theta|\approx N\Big(4d^{2}+(2\times 4)d^{2}\Big)=12Nd^{2}$ (4)
@@ 3.2 Improved Flexibility through Mixture of LoRAs
A standard LoRA module replaces a weight matrix $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ with a low-rank update:
$W^{\prime}=W+\frac{\alpha}{r}BA$ (9)
where $B\in\mathbb{R}^{d_{\text{out}}\times r}$ and $A\in\mathbb{R}^{r\times d_{\text{in}}}$ are trainable matrices of rank $r\ll d_{\text{in}},d_{\text{out}}$, drastically reducing the number of additional parameters.
@@
$\text{FFN}(h;\theta,\psi_{i})=\\ \big(W_{\mathrm{up}}+\tfrac{\alpha}{r}B_{i,2}A_{i,2}\big)\,\sigma\!\Big(\big(W_{\mathrm{down}}+\tfrac{\alpha}{r}B_{i,1}A_{i,1}\big)h\Big)$ (10)
@@ 3.5 ModernALBERT Variants
We design four variants of ModernALBERT to target different model scales:
Tiny 14 7 6, 7 768 1152 / 2624
Medium 12 3 3 1024 2624 / 4096
Base 16 4 3,4 1024 2624 / 4096
Large 24 6 3,4,5,6 1024 2624 / 4096
This design balances model depth, parameter efficiency, and representational flexibility across different scales, enabling ModernALBERT to achieve high performance while maintaining compactness and efficiency.
4 Experimental Results
Model F1 (%) Exact (%)
BERT-base 88.6 80.6
RoBERTa-base 91.7 84.7
ALBERT-xxlarge 92.5 84.5
ModernBERT-base 92.6 85.2
Ours
ModernALBERT-tiny 90.0 82.9
ModernALBERT-medium 90.4 82.9
ModernALBERT-base 92.8 86.1
ModernALBERT-large 92.9 85.9
Table 2: Results on the SQuAD-v2 Benchmark.
We evaluate ModernALBERT on natural language understanding, question answering, and information retrieval benchmarks. Our goal is to show that adding the Mixture-of-LoRAs (MoL) layer to a recursive, parameter-shared architecture provides substantial efficiency gains without sacrificing representational capacity. As shown in Table 3, ModernALBERT performs strongly across model scales. The largest variant, ModernALBERT-large (120M), achieves a GLUE average of 88.72, surpassing the fully parameterised ModernBERT-base (149M, 88.45) and outperforming previous compact state-of-the-art models such as NomicBERT and MosaicBERT. Smaller variants, including ModernALBERT-base (75M), remain competitive with models nearly twice their size, validating the effectiveness of conditional computation in recursive architectures.
4.1 GLUE Benchmark Results
@@ 5 Optimising MoL for Inference
The proposed MoL layer extends a parameter-shared recursive transformer by introducing conditional computation. Each MoL layer consists of a shared feed-forward backbone augmented with several lightweight LoRA modules that act as experts, coordinated by a small routing network. This design captures the expressive flexibility of Mixture-of-Experts architectures while preserving the compactness of parameter sharing. However, conditional routing requires evaluating multiple experts per token, introducing latency and memory overhead at inference time. Understanding this trade-off is key to integrating MoL-based models in practical deployments.
5.1 Representational Power of Expert Specialisation
@@
The increase in capacity in the MoL architecture arises from expert specialisation. As the router learns to activate different LoRA modules for different inputs, each module becomes responsible for modelling particular aspects of the data distribution (e.g., syntactic cues, semantic relations, or domain-specific patterns). This division of responsibility yields richer internal representations and helps mitigate the expressivity limitations inherent to fully shared architectures. Although the additional experts provide performance gains, dynamic routing increases inference cost because several experts may be activated per token. (shazeer2017mixture; lepikhin2020gshard; fedus2021switch)
5.2 Addressing Computational Complexity through Expert Merging
To reduce inference costs, we replace runtime routing with a single static adapter obtained by merging expert parameters (he2023mergingexpertsoneimproving). This consolidation eliminates conditional branching during inference and enables computation through a single dense update, producing substantial latency reductions. As shown in Table 7, ModernALBERT variants using merged experts maintain favourable efficiency characteristics. For instance, ModernALBERT-tiny achieves 9.46 ms latency, 106,527 tokens/s throughput, and only 0.196 GB of GPU memory usage—substantially outperforming ModernBERT-base.
We study two representative merging strategies.
5.2.1 Uniform Averaging
The simplest method assigns equal weight to all experts:
$w_{i}=\frac{1}{E}\quad\text{for all }i=1,\dots,E.$ (13)
This procedure is computationally trivial and discards token-level routing dynamics, yet it produces strong results. In our experiments (Table 8), uniform averaging preserves most of the representational quality of the unmerged model, with minimal degradation across MNLI, STS-B, RTE, and SST-2.
5.2.2 Exponentially Weighted Merging
A more informed merging strategy retains coarse routing information without per-token computation by maintaining an exponential moving average (EMA) of the router’s activation patterns. Let $p_{i,t}\in\mathbb{R}^{E}$ be the router probability vector over $E$ experts for token $t$ in sample $i$. We first compute a per-sample routing vector by averaging over tokens:
$r_{i}=\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}p_{i,t},$ (14)
where $T_{i}$ is the number of non-padding tokens. The batch-level routing vector is then obtained by averaging over the batch:
$r_{b}=\frac{1}{B}\sum_{i=1}^{B}r_{i}.$ (15)
This $r_{b}$ serves as a single EMA “sample” to update the global routing weights:
$w_{t}=\alpha w_{t-1}+(1-\alpha)r_{b},$ (16)
with normalisation to ensure $\sum_{j}w_{t,j}=1$. The resulting $w_{t}$ can be used to merge expert parameter deltas into a single adapter:
$\Delta=\sum_{j=1}^{E}w_{t,j}\,\Delta_{j}.$ (17)
In our experiments, we found that vanilla uniform averaging generally preserves accuracy slightly better, though EMA merging can be beneficial in scenarios with more heterogeneous or domain-specific routing patterns.
Method MNLI STS-B RTE SST-2
No Merging 88.2 91.8 85.6 95.2
Vanilla 87.9 91.5 84.83 94.8
EMA 87.8 91.4 83.75 94.7
Table 8: Effect of Expert Merging on Different Datasets.
Overall, our experiments indicate that vanilla uniform averaging is the most reliable merging strategy in practice. As shown in Table 8, it yields slightly stronger accuracy preservation than EMA across MNLI, STS-B, and SST-2, while maintaining a minimal computational profile. We note, however, that EMA-based merging remains valuable in settings where expert activation patterns are more heterogeneous or domain-specific, as it captures longer-term routing trends that uniform averaging ignores. In such cases, leveraging historical routing statistics may provide better generalisation, especially when models are exposed to non-stationary or task-adaptive input distributions.

### v2

@@ Abstract
Parameter sharing in recursive transformers reduces model size but collapses layer-wise expressivity. We propose Mixture of LoRAs (MoL), a lightweight mechanism that uses low-rank experts to simulate a Mixture-of-Experts layer. By replacing feed-forward networks (FFNs) in the shared transformer with these conditional modules, we effectively restore the expressivity lost due to parameter sharing. We pretrain a modernised recursive architecture, ModernALBERT, integrating rotary embeddings, GeGLU, FlashAttention, and a distillation-based initialisation. Across GLUE, SQuAD-v2, and BEIR, ModernALBERT (50M–120M) achieves state-of-the-art performance among compact models and surpasses larger fully parameterised baselines. We also propose an expert-merging procedure that compresses MoL into a single adapter at inference while preserving accuracy, enabling efficient deployment. Our results show that conditional computation effectively restores the expressivity lost under aggressive parameter sharing in recursive transformers.  [^1] Our models are available at https://huggingface.co/collections/nlpie/modern-recursive-transformers
@@ 1 Introduction
Existing approaches attempt to mitigate the expressivity loss caused by parameter sharing in three main ways. (i) Static adapters (e.g., bottleneck adapters or LoRAs) add depth-specific parameters but do not adapt at the token level and do not modify the shared FFN itself. (ii) Mixture-of-Adapters (MoA) and LoRA-based MoE variants introduce dynamic experts after the FFN, leaving its internal transformation unchanged. (iii) Relaxed Recursive Transformers (RRT) attach static LoRA modules to a shared block, partially restoring depth-specific behaviour but lacking token-dependent routing. Crucially, none of these designs introduce conditional computation inside the shared FFN during pretraining. In this work, we explore whether such internal modulation can more effectively restore the expressivity lost due to parameter sharing.
To address these limitations, we introduce ModernALBERT, a recursive transformer that restores expressivity through Mixture of LoRAs (MoL), a mechanism for conditional computation within the shared architecture. We replace select FFNs in the shared transformer with MoL modules, embedding a small pool of LoRA experts directly into the network. By routing tokens to a sparse subset of these experts, MoL allows the shared backbone to exhibit distinct behaviours for different inputs. This approach reestablishes the layer-wise functional diversity that parameter sharing typically eliminates, all while preserving strict parameter efficiency (hu2021lora).
@@
  • We introduce Mixture of LoRAs (MoL), a conditional-computation layer that integrates low-rank experts into a shared FFN, restoring expressivity in recursive transformers.
@@ 3.1 Parameter-Sharing through Recursions
$|\Theta|\approx N\Big(4d^{2}+2\times(4d^{2})\Big)=12Nd^{2}$ (4)
@@ 3.2 Improved Flexibility through Mixture of LoRAs
A standard LoRA module replaces a weight matrix $W\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ with a low-rank update:
$W^{\prime}=W+\frac{\alpha}{r}AB$ (9)
where $A\in\mathbb{R}^{d_{\text{in}}\times r}$ and $B\in\mathbb{R}^{r\times d_{\text{out}}}$ are trainable matrices of rank $r\ll d_{\text{in}},d_{\text{out}}$, drastically reducing the number of additional parameters.
@@
$\text{FFN}(h;\theta,\psi_{i})=\\ \big(W_{\mathrm{up}}+\tfrac{\alpha}{r}A_{i,2}B_{i,2}\big)\,\sigma\!\Big(\big(W_{\mathrm{down}}+\tfrac{\alpha}{r}A_{i,1}B_{i,1}\big)h\Big)$ (10)
@@ 3.5 ModernALBERT Variants
We design four variants of ModernALBERT (Table 1) to target different model scales. This design balances model depth, parameter efficiency, and representational flexibility across different scales, enabling ModernALBERT to achieve high performance while maintaining compactness and efficiency.
@@
4 Experimental Results
Model F1 (%) Exact (%)
BERT-base 88.6 80.6
RoBERTa-base 91.7 84.7
ALBERT-xxlarge 92.5 84.5
ModernBERT-base 92.6 85.2
Ours
ModernALBERT-tiny 90.0 82.9
ModernALBERT-medium 90.4 82.9
ModernALBERT-base 92.8 86.1
ModernALBERT-large 92.9 85.9
Table 3: Results on the SQuAD-v2 Benchmark.
We evaluate ModernALBERT on natural language understanding, question answering, and information retrieval benchmarks. Our goal is to show that adding the Mixture-of-LoRAs (MoL) layer to a recursive, parameter-shared architecture provides substantial efficiency gains without sacrificing representational capacity. As shown in Table 2, ModernALBERT performs strongly across model scales.
4.1 GLUE Benchmark Results
@@ 5 Optimising MoL for Inference
The proposed MoL layer augments a parameter-shared recursive transformer with lightweight LoRA experts coordinated by a routing network. While this design captures the flexibility of Mixture-of-Experts architectures, the conditional routing mechanism introduces latency overhead during inference by evaluating multiple experts per token. To address this, we propose compressing the MoL layer into a single static adapter for deployment.
@@
5.1 Expert Merging Strategies
To eliminate conditional branching, we synthesise a single dense adapter $\Delta_{\text{merged}}$ from the pool of $E$ experts $\{\Delta_{1},\dots,\Delta_{E}\}$. This aggregation is defined by a weighting vector $\mathbf{w}\in\mathbb{R}^{E}$:
$\Delta_{\text{merged}}=\sum_{j=1}^{E}w_{j}\Delta_{j}.$ (13)
During fine-tuning, the dynamic routing mechanism is disabled, and the model updates $\Delta_{\text{merged}}$ directly (he2023mergingexpertsoneimproving). This procedure effectively converts the sparse MoL layer into a standard, static LoRA module, reducing the inference cost to that of a dense model. As shown in Table 7, this approach yields substantial efficiency gains; ModernALBERT-tiny achieves 9.46 ms latency and 106,527 tokens/s throughput while using only 0.196 GB of GPU memory.
We investigate two strategies for managing this aggregation during fine-tuning.
5.1.1 Uniform Initialization
In the first strategy, we initialise the merged adapter by averaging the pre-trained experts uniformly ($w_{j}=1/E$). At each step, the merged adapter $\Delta_{\text{merged}}$ is formed by averaging the experts, and the model parameters are updated via standard backpropagation. This method assumes a neutral starting point, allowing the aggregate representation to adapt to the downstream task without reliance on token-specific routing.
5.1.2 Dynamic EMA Merging
The second strategy updates the merging weights $\mathbf{w}$ dynamically throughout fine-tuning based on the router’s activation patterns. We initialise $\mathbf{w}$ uniformly ($w_{j}=1/E$), allowing the EMA updates to gradually refine the distribution from a neutral starting point. Let $p_{i,t}\in\mathbb{R}^{E}$ be the router probability vector for token $t$ in sample $i$. We first compute a per-sample routing vector $r_{i}$ and a batch-level average $r_{b}$:
$r_{i}=\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}p_{i,t},\quad\quad r_{b}=\frac{1}{B}\sum_{i=1}^{B}r_{i}.$ (14)
We then update the global merging weights $\mathbf{w}$ using an exponential moving average (EMA) of these batch statistics:
$\mathbf{w}_{\text{new}}=\alpha\mathbf{w}_{\text{old}}+(1-\alpha)r_{b},$ (15)
where $\alpha$ is a decay factor. Regarding the router itself, we adopt a task-dependent approach: for data-rich tasks like MNLI, the router is trained jointly to refine expert selection; for smaller datasets (STS-B, RTE, SST-2), the router is frozen to leverage pre-trained specialisation while only the EMA weights adapt.
As shown in Table 8, the EMA-based strategy consistently outperforms uniform initialization (Vanilla). By dynamically adjusting the expert contributions based on routing statistics, EMA preserves nearly all of the unmerged model’s accuracy, with particularly strong gains on RTE (86.28 vs 84.83) and SST-2. These results confirm that incorporating the router’s learned preferences into the merging process, even when the router itself is frozen, provides a superior adaptation mechanism compared to a static uniform initialization.
6 Discussion
@@
The increase in capacity in the MoL architecture arises primarily from expert specialisation. As the router learns to activate different LoRA modules for different inputs, each module becomes responsible for modelling particular aspects of the data distribution, such as syntactic cues, semantic relations, or domain-specific patterns. This division of responsibility yields richer internal representations and helps mitigate the expressivity limitations inherent to fully shared architectures. Although the additional experts provide clear performance gains, dynamic routing naturally increases inference cost because several experts may be activated per token (shazeer2017mixture; lepikhin2020gshard; fedus2021switch).
Our results on expert merging further suggest that while dynamic routing is crucial for training high-capacity models, the learned expressivity can often be compressed into a static representation for inference. The success of the EMA strategy implies that the aggregate behaviour of the experts captures a robust policy that adapts well across general domains. This finding aligns with recent work on model merging and distillation, indicating that the computational cost of conditional branching is beneficial for learning but not always strictly necessary for maintaining performance during deployment. This trade-off allows MoL-based models to occupy a unique position where they train with the expressivity of large MoEs but deploy with the speed and footprint of compact dense models.
@@ 8 Conclusion and Future Work
Method MNLI STS-B RTE SST-2
No Merging 88.2 91.8 85.6 95.2
Vanilla 87.9 91.5 84.83 94.8
EMA 88.0 91.7 86.28 95.1
Table 8: Effect of Expert Merging on Different Datasets.